hartnn (hartnn):

∫[(x + 1)/(x*(1 + x*e^x)²)]dx

hartnn (hartnn):

\(\huge \int \frac{x+1}{x(1+xe^x)^2}dx=..?\)

hartnn (hartnn):

partial fractions? or x= log y substitution or how ??

OpenStudy (abb0t):

Partial fractions.

hartnn (hartnn):

for partial fractions, there should be linear expression, 1+xe^x is not linear....

hartnn (hartnn):

*polynomial expression

OpenStudy (abb0t):

But you do have a function squared. wHICH You can expand to get a ^2x

hartnn (hartnn):

what would i take in numerator ?? if it were (x+1)^2, i would have taken A/(x+1) +B/(x+1)^2 ...

OpenStudy (abb0t):

Yeah, or Ax+B for numerator if you expand it.

hartnn (hartnn):

yeah, what here ?

OpenStudy (abb0t):

I think you can use Ax+B w/o expanding. haha. I'm just making guesses now. I don't have a pen within a 12 inch radius Lol

OpenStudy (abb0t):

But it looks like partial should work.

hartnn (hartnn):

the answer suggested me partial, but how is the question....

OpenStudy (abb0t):

\[\frac{ A}{ x } + \frac{ Bx+C }{ (1+xe^x) }+...\]

hartnn (hartnn):

not justified, nor will give complete answer...

OpenStudy (abb0t):

hwat?! :P

OpenStudy (anonymous):

Right but what are the steps to derive this?

OpenStudy (abb0t):

~M*A*G*I*C~

hartnn (hartnn):

got it :)

hartnn (hartnn):

put 1+xe^x = t

hartnn (hartnn):

then partial fractions...

OpenStudy (abb0t):

nicely done, mate :)

hartnn (hartnn):

\(\huge \int \frac{e^x(x+1)}{x.e^x(1+xe^x)^2}dx=\int \frac{dy}{y^2(y-1)}\)

OpenStudy (anonymous):

It's not magic anymore .. :(

OpenStudy (mathmate):

:)